OpenStudy (anonymous):

integrate: 0 to i:nfinity ʃ e^tx 1/6(e^-x/2 +xe^-x/2

OpenStudy (kainui):

\[\int\limits_{\infty}^{0}e^{x}/(6e^{-x/2}+xe^{-x/2})dx\] Is that what you're trying to integrate?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\frac{1}{6} e^t x \left(\frac{e^{-x}}{2}+\frac{e^{-x} x}{2}\right) \]Is this the problem expression?

OpenStudy (anonymous):

no can i draw it for u?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):